OpenStudy (dls):

How much mass is converted into energy per day in a power plant operated at 10^6kW

OpenStudy (dls):

@Mashy @yrelhan4

OpenStudy (dls):

\[\LARGE P=\frac{w}{T}\] I used this to get work done=energy

OpenStudy (dls):

T=24*60*60

OpenStudy (dls):

then used \[\LARGE E=mc^{2}\]

OpenStudy (anonymous):

seems legit

OpenStudy (dls):

am I right anyhow

OpenStudy (dls):

WOAH :O but the answer i got is weird :/

OpenStudy (yrelhan4):

chaa gaya puttar. :P

OpenStudy (dls):

\[5.4 \times 10^{-7} KW/m^{2}{s}^{-4}\]

OpenStudy (dls):

weird units :O

OpenStudy (anonymous):

what!? .. ofcourse.. conver W to j per second!

OpenStudy (dls):

i didnt know how to do that :P was unsure so left it like that

OpenStudy (dls):

multiplying by 1.6*10^19 but unsure

OpenStudy (dls):

is it wrong if I left it in that conditon :/

OpenStudy (anonymous):

what?! :O.. and besides m getting answer as 288kg

OpenStudy (dls):

wtf :|

OpenStudy (dls):

method is right?

OpenStudy (anonymous):

oh wait.. i forgot to do c square :D

OpenStudy (dls):

-_-

OpenStudy (dls):

@yrelhan4 im scared of his reply :// stop him!!

OpenStudy (dls):

he will prove me wrong like everytime..:'(

OpenStudy (yrelhan4):

go mashy go. :D

OpenStudy (anonymous):

but still m getting different answer..!! \[E = Pt = mc^2 => m = Pt/c^2 => \frac{ 10^9 \times 24 \times 3600}{ 9 \times 10^{18}}\]

OpenStudy (dls):

I didn't change kW to W or joules was quite unsure is it wrong like that? o.O

OpenStudy (anonymous):

you have to put everything in S.I so you need to put it in Watts.. Watts or J/s is one and the same thing anyways.. how did you get the number wrong? :P

OpenStudy (dls):

i guess i took speed of light as 3.8 * 10^8 LOL..

OpenStudy (dls):

yeah i took it 3.8 into 10^8 -_-

OpenStudy (dls):

is it wrong..........? :P

OpenStudy (anonymous):

no no.. its your wish .. you can take whatever you want to be the speed of light

OpenStudy (dls):

it doesnt really matter you know :P i approxed 3.8 into 4

OpenStudy (dls):

4*10^8 !!

OpenStudy (anonymous):

but then you ll have to redefine the meter.. thats all!

OpenStudy (dls):

my method is right :D thats all matters :P

OpenStudy (anonymous):

its 3 you fool.. not 4 ! :O ..

OpenStudy (dls):

i took it as 3.8 then approxed that to 4 :P

OpenStudy (anonymous):

HOW CAN YOU DO THAT?! :O.. its not 3.8 !!!! its 3!

OpenStudy (dls):

i know :| typo in exam :P

OpenStudy (anonymous):

that was in your exam? :O

OpenStudy (dls):

yes

OpenStudy (dls):

aaj tha